# Real-to-Sim via End-to-End Differentiable Simulation and Rendering

### v1

@@ VI Limitations
Our method assumes ground-truth object masks from the scene, which might not always be possible even with advanced foundational segmentation models. In addition, our method does not consider more advanced appearance models, lighting sources, and shadows. As a result, the rendered scenes could have artifacts that do not match the real-world observations. Currently, each optimization run is completed in under 15 mins on a standard PC with an Intel i9-13900KF CPU, 64 GB of RAM, and a GeForce RTX 4090 GPU. While this is not ideal for online robotics applications, we intend to reduce the runtime by using better initial guesses of geometry and physical parameters from data-driven pre-trained models and more efficient implementation.

### v3

@@ VI Limitations
Our method assumes ground-truth object masks from the scene, which might not always be possible even with advanced foundational segmentation models. In addition, our method does not consider more advanced appearance models, lighting sources, and shadows. As a result, the rendered scenes could have artifacts that do not match the real-world observations. Currently, each optimization run is completed in under 15 mins on a standard PC with an Intel i9-13900KF CPU, 64 GB of RAM, and a GeForce RTX 4090 GPU. While this is not ideal for online robotics applications, we intend to reduce the runtime by using better initial guesses of geometry and physical parameters from data-driven pre-trained models and more efficient implementation. Finally, our method currently struggles on objects whose rotation can not be properly identified from a surface point cloud, such as a cylinder. We aim to explore tracking methods that also leverages surface textures for pose tracking in future work.